# A Nonconforming Virtual Element Method for Advection-Diffusion-Reaction Problems with CIP Stabilization

### v1

@@ 3.2 Inf-Sup condition
The goal of this section is to prove the following inf-sup condition for the discrete problem (26):
@@
$+\dfrac{\varepsilon}{\delta h}\|{\Pi^{0,\partial E}_{k-1}}v_{h}\|^{2}_{0,\partial E\cap\Gamma}+\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\partial E\cap\Gamma_{\text{in}}}+J_{h}^{E}(v_{h},v_{h})\,.$
@@
as the second argument in $\mathcal{A}_{\rm cip}(\cdot,\cdot)$. This is not possible, since $\widetilde{w}_{h}\not\in V^{k,nc}_{h}(\Omega_{h})$. To avoid this problem, we take its Oswald interpolant $\pi\widetilde{w}_{h}$ instead. Then, the difference between $\widetilde{w}_{h}$ and $\pi\widetilde{w}_{h}$ is controlled thanks to the jump term. We combine these two results and conclude the proof of the inf-sup condition (56) in Theorem 3.2.
@@
Under assumptions (A1), given $v_{h}\in V_{h}(\Omega_{h})$, we have
$\begin{split}\mathcal{A}_{\rm cip}(v_{h},v_{h})\gtrsim&\varepsilon\|\nabla_{\!h}v_{h}\|_{0}^{2}+\sigma\|v_{h}\|_{0}^{2}+\sum_{E\in\Omega_{h}}\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}\\ &+\sum_{E\in\Omega_{h}}\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|_{0,\partial E\cap\Gamma_{{\rm in}}}^{2}+J_{h}(v_{h},v_{h})\,.\end{split}$
@@
$\leq 2\varepsilon\delta h_{E}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}^{E}\|_{\partial E\cap\Gamma}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{e}^{2}$
$\leq 2\varepsilon\delta C_{\rm tr}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\|_{0,E}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{e}^{2}$
$\leq 2\varepsilon\delta C_{\rm tr}\|\nabla v_{h}\|_{0,E}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{e}^{2}\,,$
@@
gives the result.
{comment}
To handle the last non-symmetric term, we note that
$2\langle\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n},v_{h}\rangle_{\Gamma_{E}}=2\langle\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n},{\Pi^{0,\partial E}_{k-1}}v_{h}\rangle_{\Gamma_{E}}\,.$
Using Cauchy-Schwarz inequality and Young inequality for a constant $\alpha>0$
$\varepsilon\|\nabla v_{h}\|^{2}_{0,E}-2\varepsilon\langle\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n}^{E},{\Pi^{0,\partial E}_{k-1}}v_{h}\rangle_{\Gamma_{E}}+\dfrac{\varepsilon}{\delta h}\|{\Pi^{0,\partial E}_{k-1}}v_{h}\|_{0,\Gamma_{E}}^{2}\\ \geq\varepsilon\|\nabla v_{h}\|^{2}_{0,E}-\dfrac{h\,\varepsilon}{\alpha}\|\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n}^{E}\|_{0,\Gamma_{E}}^{2}+\left(\dfrac{1}{\delta}-\dfrac{\alpha}{2}\right)\dfrac{\varepsilon}{h}\|{\Pi^{0,\partial E}_{k-1}}v_{h}\|^{2}_{0,\Gamma_{E}}\,.$
Under assumptions (A1), on the second term, we can use the polynomial trace inequality, to obtain
$h\|\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n}^{E}\|_{0,\Gamma_{E}}^{2}\leq C_{t}\|\nabla{\Pi^{\nabla,E}_{k}}v_{h}\|_{0,E}^{2}\leq C_{t}\|\nabla v_{h}\|^{2}_{0,E}\,,$
for a uniform constant $C_{t}>0$. Setting $\alpha=C_{t}$ and $0<\delta<C_{t}$, we obtain
$\varepsilon\|\nabla v_{h}\|^{2}_{0,E}-2\varepsilon\langle\nabla{\Pi^{\nabla,E}_{k}}v_{h}\cdot\boldsymbol{n}^{E},{\Pi^{0,\partial E}_{k-1}}v_{h}\rangle_{\Gamma_{E}}+\dfrac{\varepsilon}{\delta h}\|v_{h}\|_{0,\Gamma_{E}}^{2}\gtrsim\varepsilon\|v_{h}\|^{2}_{0,E}+\dfrac{\varepsilon}{\delta h}\|{\Pi^{0,\partial E}_{k-1}}v_{h}\|^{2}_{0,\Gamma_{E}}\,.$
Combing all of these results and summing over all the polygons, we obtain
$\varepsilon\|\nabla v_{h}\|_{0}^{2}+J_{h}(v_{h},v_{h})+\sigma\|v_{h}\|_{0}^{2}+\|{\Pi^{0,\partial E}_{k-1}}v_{h}\|^{2}_{0,\Gamma}+\|{\Pi^{0,E}_{k}}v_{h}\|_{0,\Gamma}^{2}\lesssim\mathcal{A}_{\rm cip}(v_{h},v_{h})\,.$
@@
where $\boldsymbol{\beta}_{h}$ is the $L^{2}$-projection of $\boldsymbol{\beta}$ in the space of piecewise linear functions $\mathbb{P}_{1}(\Omega_{h})$. Then, under assumptions (A1), if the mesh size satisfies $h>\varepsilon$, we have
@@
We remark that $\mathcal{N}^{E}_{h}(\cdot,\cdot)$ is different from zero only for $E$ with at least one edge on $\Gamma$.
@@
For $\eta_{\mathcal{N}_{2}}$, we use the orthogonality of ${\Pi^{0,\partial E}_{k-1}}$, the Cauchy-Schwarz inequality, inverse trace and inverse inequalities, and $\varepsilon<h$, to obtain
@@
For $\eta_{\mathcal{N}_{3}}$, thanks to orthogonality, we remove the ${\Pi^{0,\partial E}_{k-1}}$ projection on the second term and we proceed similarly to the previous cases:
@@
$\geq h\,\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}-C\,h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{W^{1,\infty}(E)}h\|\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\geq h\,\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}-C\,h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0,E}\,,$
where in the third step we have used $\|\boldsymbol{\beta}_{h}-\boldsymbol{\beta}\|_{L^{\infty}(E)}\lesssim h|\boldsymbol{\beta}|_{W^{1,\infty}(E)}$.
@@
where we used that $\gamma^{2}_{e}\leq\gamma_{e}$ (since $\gamma_{e}\leq 1$, see (7)). On each $e$, the argument in the first sum on the right-hand side of the previous inequality is controlled using the trace inequality and standard estimates on $\boldsymbol{\beta}\in W^{1,\infty}(\Omega)$:
$h^{2}\|[\![(\boldsymbol{\beta}_{h}-\boldsymbol{\beta})\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}]\!]\|_{0,e}^{2}$ $\lesssim h^{4}|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(E\cup E^{\prime})}h^{-1}\|\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|_{0,E\cup E^{\prime}}^{2}$ (77)
$\lesssim h|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(E\cup E^{\prime})}\|{\Pi^{0}_{k}}v_{h}\|_{0,E\cup E^{\prime}}^{2}$
$\lesssim h|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(E\cup E^{\prime})}\|v_{h}\|_{0,E\cup E^{\prime}}^{2}\,,$
@@
$-C\Big{(}h|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(\mathcal{D}(E))}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}+\sum_{E^{\prime}\subset\mathcal{D}(E)}J_{h}^{E^{\prime}}(v_{h},v_{h})\Big{)}\,.$
It remains to control $\eta_{\boldsymbol{\beta}_{3}}$. Since $\boldsymbol{\beta}_{h}\in\mathbb{P}_{1}(E)$, it holds $\bigl{(}\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h},({\Pi^{0,E}_{k}}-I)w_{h}\bigr{)}_{0,E}=0$. Hence we have
@@
$\gtrsim-|\boldsymbol{\beta}|_{W^{1,\infty}(E)}h\|\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h\|\boldsymbol{\beta}_{h}\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|_{0,\mathcal{D}(E)}$
$\gtrsim-|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0.\mathcal{D}(E)}^{2}$
@@
$+h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0,E}$
$+h|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(\mathcal{D}(E))}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}+|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}\Big{)}\,.$
@@
$+J_{h}(v_{h},v_{h})+\sum_{E\in\Omega_{h}}\left(h|\boldsymbol{\beta}|^{2}_{W^{1,\infty}(E)}+|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\right)\|v_{h}\|_{0,E}^{2}$
$+\sum_{E\in\Omega_{h}}h^{\frac{1}{2}}\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0,E}\Big{)}\,.$
Above, we have also used the property that, due to assumption (A1), summing over the elements, each polygon is counted only a uniformly bounded number of times, even when the terms involve norms on $\mathcal{D}(E)$ or $\mathcal{D}(\mathcal{D}(E))$.
@@
$h^{\frac{1}{2}}\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\lesssim h^{\frac{1}{2}}\left(\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}+|\boldsymbol{\beta}|_{W^{1,\infty}(E)}\|v_{h}\|_{0,E}\right).$ (83)
@@
We also need the following estimate.
@@
$\dfrac{\varepsilon}{\delta h}\|{\Pi^{0,\partial E}_{k-1}}w_{h}\|^{2}_{0,\partial E\cap\Gamma}\leq\dfrac{\varepsilon}{\delta h}\|w_{h}\|^{2}_{0,\partial E\cap\Gamma}\lesssim\varepsilon\|\pi(\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h})\|^{2}_{0,E}\lesssim h\|\boldsymbol{\beta}_{h}\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}\,,$ (87)
@@
{comment}
Recalling the definition of $w_{h}$, we easily have that
$\varepsilon\|\nabla w_{h}\|^{2}_{0}\lesssim\varepsilon\|\pi\boldsymbol{\beta}_{h}\cdot{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}\lesssim\varepsilon\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}\lesssim h\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}\,.$ (90)
Using (83), we conclude
$\varepsilon\|\nabla w_{h}\|^{2}_{0}\lesssim h\Bigl{(}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}+|\boldsymbol{\beta}|_{W^{1,\infty}(E)}^{2}\|v_{h}\|^{2}_{0,\mathcal{D}(E)}\Bigr{)}\,.$ (91)
On the other hand, we have that
$\sigma\|w_{h}\|^{2}_{0,E}\lesssim\sigma\|h\pi\boldsymbol{\beta}_{h}\cdot{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\lesssim\sigma\|v_{h}\|^{2}_{0,\mathcal{D}(E)}\,.$ (92)
The convective term in the norm is controlled by
$h\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}w_{h}\|_{0,E}^{2}\lesssim h^{-1}\|{\Pi^{0,E}_{k}}w_{h}\|_{0,E}^{2}\lesssim h\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,\mathcal{D}(E)}^{2}\,,$ (93)
and using again (83). The two boundary term are controlled as
$\dfrac{\varepsilon}{\delta h}\|{\Pi^{0,\partial E}_{k-1}}w_{h}\|^{2}_{0,\Gamma_{E}}\leq\dfrac{\varepsilon}{\delta h}\|w_{h}\|^{2}_{0,\Gamma_{E}}\lesssim\varepsilon\|\pi\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}\leq h\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}\,,$ (94)
and
$\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}w_{h}\|^{2}_{0,\Gamma_{E}}\lesssim h^{-1}\|{\Pi^{0,E}_{k}}w_{h}\|^{2}_{0,E}\lesssim h\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,\mathcal{D}(E)}^{2}\,,$ (95)
and then using again (83)  . The proof is completed by gathering (91), (92), (93), (94), (95) and (65) (in combination with (83)).
If $h\lesssim\varepsilon$, from Lemma 3.3, the definition of the norm $\|\cdot\|_{\rm cip}$, and the estimate

### v2

@@ 3.2 Inf-Sup condition
The goal of this section is to prove the following inf-sup condition for the discrete problem (27), namely
@@
$+\dfrac{\varepsilon}{\delta h}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}+\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,\partial E\cap\Gamma_{\text{in}}}+J_{h}^{E}(v_{h},v_{h})\,.$
@@
as the second argument in $\mathcal{A}_{\rm cip}(\cdot,\cdot)$. This is not possible, since $\widetilde{w}_{h}\not\in V^{k,nc}_{h}(\Omega_{h})$, and we take its Oswald interpolant $\pi\widetilde{w}_{h}$ instead. Then, the difference between $\widetilde{w}_{h}$ and $\pi\widetilde{w}_{h}$ is controlled thanks to the jump term. We combine these two results and conclude the proof of the inf-sup condition (51) in Theorem 3.2.
@@
Under assumptions (A1), given $v_{h}\in V^{k,nc}_{h}(\Omega_{h})$, we have
$\begin{split}\mathcal{A}_{\rm cip}(v_{h},v_{h})\gtrsim&\varepsilon\|\nabla_{\!h}v_{h}\|_{0}^{2}+\sigma\|v_{h}\|_{0}^{2}+\sum_{E\in\Omega_{h}}\dfrac{\varepsilon}{\delta h}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}\\ &+\sum_{E\in\Omega_{h}}\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|_{0,\partial E\cap\Gamma_{{\rm in}}}^{2}+J_{h}(v_{h},v_{h})\,.\end{split}$
@@
$\leq 2\varepsilon\delta h_{E}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}^{E}\|_{\partial E\cap\Gamma}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}$
$\leq 2\varepsilon\delta C_{\rm tr}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\|_{0,E}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}$
$\leq 2\varepsilon\delta C_{\rm tr}\|\nabla v_{h}\|_{0,E}^{2}+\dfrac{\varepsilon}{2\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi_{k-1}^{0,e}v_{h}\|_{0,e}^{2}\,,$
@@
gives the result. ∎
@@
where $\boldsymbol{\beta}_{h}$ is the $L^{2}$-projection of $\boldsymbol{\beta}$ in the space of piecewise linear functions $[\mathbb{P}_{1}(\Omega_{h})]^{d}$. Then, under assumptions (A1), if the mesh size satisfies $h>\varepsilon$, we have
@@
We remark that $\mathcal{N}^{E}_{h}(\cdot,\cdot)$ is different from zero only if $E$ has at least one facet on $\Gamma$.
@@
For $\eta_{\mathcal{N}_{2}}$, we use the orthogonality of $\Pi^{0,e}_{k-1}$, the Cauchy-Schwarz inequality, inverse trace and inverse inequalities, and $\varepsilon<h$, to obtain
@@
For $\eta_{\mathcal{N}_{3}}$, thanks to orthogonality, we remove the $\Pi^{0,e}_{k-1}$ projection on the second term and we proceed similarly to the previous cases:
@@
$\geq h\,\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}-C\,h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}h\|\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\geq h\,\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|^{2}_{0,E}-C\,h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0,E}\,,$
where in the third step we have used $\|\boldsymbol{\beta}_{h}-\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}\lesssim h|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}$.
@@
where we used that $\gamma^{2}_{e}\leq\gamma_{e}$ (since $\gamma_{e}\leq 1$, see (7)). On each $e$, the argument in the first sum on the right-hand side of the previous inequality is controlled using the trace inequality and standard estimates on $\boldsymbol{\beta}\in[W^{1,\infty}(\Omega)]^{d}$:
$h^{2}\|[\![(\boldsymbol{\beta}_{h}-\boldsymbol{\beta})\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}]\!]\|_{0,e}^{2}$ $\lesssim h^{4}|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(E\cup E^{\prime})]^{d}}h^{-1}\|\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|_{0,E\cup E^{\prime}}^{2}$ (72)
$\lesssim h|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(E\cup E^{\prime})]^{d}}\|{\Pi^{0}_{k}}v_{h}\|_{0,E\cup E^{\prime}}^{2}$
$\lesssim h|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(E\cup E^{\prime})]^{d}}\|v_{h}\|_{0,E\cup E^{\prime}}^{2}\,,$
@@
$-C\Big{(}h|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(\mathcal{D}(E))]^{d}}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}+\sum_{E^{\prime}\subset\mathcal{D}(E)}J_{h}^{E^{\prime}}(v_{h},v_{h})\Big{)}\,.$
It remains to control $\eta_{\boldsymbol{\beta}_{3}}$. Since $\boldsymbol{\beta}_{h}\in[\mathbb{P}_{1}(E)]^{d}$, it holds $\bigl{(}\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h},({\Pi^{0,E}_{k}}-I)w_{h}\bigr{)}_{0,E}=0$. Hence we have
@@
$\gtrsim-|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}h\|\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h\|\boldsymbol{\beta}_{h}\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|_{0,\mathcal{D}(E)}$
$\gtrsim-|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0.\mathcal{D}(E)}^{2}\,.$
@@
$+h^{\frac{1}{2}}\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0,E}$
$+h|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(\mathcal{D}(E))]^{d}}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}+|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0,\mathcal{D}(E)}^{2}\Big{)}\,.$
@@
$+J_{h}(v_{h},v_{h})+\sum_{E\in\Omega_{h}}\left(h|\boldsymbol{\beta}|^{2}_{[W^{1,\infty}(E)]^{d}}+|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\right)\|v_{h}\|_{0,E}^{2}$
$+\sum_{E\in\Omega_{h}}h^{\frac{1}{2}}\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\,h^{\frac{1}{2}}|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0,E}\Big{)}\,.$
Above, we have also used the property that, due to assumption (A1), summing over the elements, each element is counted only a uniformly bounded number of times, even when the terms involve norms on $\mathcal{D}(E)$ or $\mathcal{D}(\mathcal{D}(E))$.
@@
$h^{\frac{1}{2}}\|\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}\lesssim h^{\frac{1}{2}}\left(\|\boldsymbol{\beta}\cdot\nabla{\Pi^{0,E}_{k}}v_{h}\|_{0,E}+|\boldsymbol{\beta}|_{[W^{1,\infty}(E)]^{d}}\|v_{h}\|_{0,E}\right).$ (78)
@@
Whenever $h>\varepsilon$, we also need the following estimate.
@@
$\sum_{e\subset\partial E\cap\Gamma}\dfrac{\varepsilon}{\delta h}\|\Pi^{0,e}_{k-1}w_{h}\|^{2}_{0,e}\leq\dfrac{\varepsilon}{\delta h}\|w_{h}\|^{2}_{0,\partial E\cap\Gamma}\lesssim\varepsilon\|\pi(\boldsymbol{\beta}_{h}\cdot\nabla{\Pi^{0,E}_{k}}v_{h})\|^{2}_{0,E}\lesssim h\|\boldsymbol{\beta}_{h}\cdot\nabla_{\!h}{\Pi^{0}_{k}}v_{h}\|^{2}_{0,\mathcal{D}(E)}\,,$ (82)
@@
If $h\leq\varepsilon$, from Lemma 3.3, the definition of the norm $\|\cdot\|_{\rm cip}$, and the estimate